# Stochastic Gravitational Wave Background from Highly-Eccentric Stellar-Mass Binaries in the Milli-hertz Band

### v1

@@ IV Discussion
Figure 4: Summary of the stochastic GW backgrounds from bursting BBHs. Here, we show the upper limit of the GW background, $\sqrt{fS_{n}(f)}$, corresponding to highly eccentric BBHs from the Globular Clusters (blue dashed lines), Galactic Field (green dashed lines), and Galactic Nuclei (brown dash-dotted lines) in the universe. Upper Panel shows the maximum level of total GW background from bursting BBHs. Bottom Panel represents the unresolvable GW noise background from bursting sources with $\rm SNR<8$ for a $4$ yr LISA mission. For demonstration purposes, we sum the contribution of the three channels mentioned above and show the collective background in the solid red lines, and plot the LISA noise curve (N2A5 configuration, in red; eLISA, in blue [88, 89]).
In Section II.2.2, we developed an analytical framework for computing the stochastic GW background from bursting sources. Subsequently, we carried out numerical simulations to determine the GW background from bursting BBHs in different formation channels, utilizing the population model of highly eccentric binaries outlined in Section III.1. Specifically, Figure 2 depicts the example of the GW stochastic background from GCs and the Galactic Fields (left and right panels, respectively), adopting a steady-state approach for the sources’ formation rate. In Figure 3, we also include the fact that a young population of stars has been observed in the center of the Milky Way [134, 144, 145, 115, 135, 116]. This young population implies that the center of our galaxy may deviate from the steady state restriction. Thus, we explore the impact of a young population in the nuclear star cluster on the stochastic background. It produces a comparable contribution to the GW background (see Figure 2, dotted green line). Additionally, there have been observational and theoretical several studies that suggested the existence of about a few thousand of solar mass inwards to S0-2’s orbit [136, 139, 140, 141, 146]. We thus add a population of compact objects inwards to S0-2’s orbit ($\sim 1000$ au). That may also have a significant contribution to the GW stochastic background.
@@
As can be seen in the Figure, highly eccentric, stellar-mass BBHs can have a non-negligible GW background contribution in the LISA detection, potentially exceeding the LISA instrumental noise at $\sim 3-7$ mHz (see Upper Panel, in Figure 4). This phenomenon is especially significant if we do not have an effective search strategy for these sources’ GW signals. In the LISA data analysis, if we only use the near-circular GW templates, the GW bursts from these sources will not be identified and will only be considered as a noise background in the detector’s output. Therefore, even with promising $\rm SNR$, the bursting GW sources will become part of the unresolvable noise background, thus affecting the LISA sensitivity (see Section II.2.1).
However, providing that all the stellar-mass bursting sources are properly identified in the future, the residual noise level is around two orders of magnitude lower than the LISA sensitivity curve, thus having a limited effect on the data analysis (see the Bottom Panel). Considering the abovementioned points, we highlight the importance of searching and identifying bursting GW sources in LISA detection. Particularly, we may need to develop portable GW templates for these sources in the matched filtering or adopt other time-frequency approaches, such as power stacking [1] and the Q-transform [93, 94], to characterize the GW bursts from astrophysical sources in LISA data analysis.
Additionally, in realistic detection, the antenna pattern of LISA undergoes rotation across the sky, causing its response to gravitational waves to change with time. As the bursting BBHs in the Milky Way predominantly contribute to the stochastic background and may have an anisotropic distribution, the amplitude of the GW burst background can vary considerably throughout the year. This observation is consistent with the pattern illustrated in Figure 1 [see, e.g., 147, for a similar case].
@@ Acknowledgements.
ZX acknowledges partial support from the Bhaumik Institute for Theoretical Physics summer fellowship. ZX, SN, and EM acknowledge the partial support from NASA ATP 80NSSC20K0505 and from NSF-AST 2206428 grant and thank Howard and Astrid Preston for their generous support. This work was supported by the Science and Technology Facilities Council Grant Number ST/W000903/1 (to BK).

### v2

@@ IV Discussion
In Section II.3, we developed an analytical framework for computing the stochastic GW background from bursting sources. Subsequently, we carried out numerical simulations to determine the GW background from bursting BBHs in different formation channels, utilizing the population model of highly eccentric binaries outlined in Section III.1. Specifically, Figure 2 depicts the example of the GW stochastic background from GCs and the Galactic Fields (left and right panels, respectively), adopting a steady-state approach for the sources’ formation rate. In Figure 3, we also include the fact that a young population of stars has been observed in the center of the Milky Way [140, 150, 151, 120, 141, 121]. This young population implies that the center of our galaxy may deviate from the steady state restriction. Thus, we explore the impact of a young population in the nuclear star cluster on the stochastic background. It produces a comparable contribution to the GW background (see Figure 2, dotted green line). Additionally, there have been observational and theoretical several studies that suggested the existence of about a few thousand of solar mass inwards to S0-2’s orbit [142, 145, 146, 147, 152]. We thus add a population of compact objects inwards to S0-2’s orbit ($\sim 1000$ au). That may also have a significant contribution to the GW stochastic background.
@@
Figure 4: Summary of the stochastic GW backgrounds from bursting BBHs. Here, we show the upper limit of the GW background, $\sqrt{fS_{n}(f)}$, corresponding to highly eccentric BBHs from the Globular Clusters (blue dashed lines), Galactic Field (green dashed lines), and Galactic Nuclei (brown dash-dotted lines) in the universe. Upper Panel shows the maximum level of total GW background from bursting BBHs. Bottom Panel represents the unresolvable GW noise background from bursting sources with $\rm SNR<8$ for a $4$ yr LISA mission. For demonstration purposes, we sum the contribution of the three channels mentioned above and show the collective background in the solid red lines, and plot the LISA noise curve (N2A5 configuration [92, 93], in orange).
@@
As can be seen in the figure, highly eccentric, stellar mass BBHs can have a non-negligible GW background contribution in the LISA detection, potentially exceeding the LISA instrumental noise at $\sim 3-7$ mHz (see Upper Panel, in Figure 4). This phenomenon is especially significant if we do not have an effective search strategy for these sources’ GW signals. In the LISA data analysis, if we only use the near-circular GW templates, the GW bursts from these sources will not be identified and will only be considered as a noise background in the detector’s output. Therefore, even with promising $\rm SNR$, the bursting GW sources will become part of the unresolvable noise background, thus affecting the LISA sensitivity (see Section II.2.1).
However, providing that all the stellar mass bursting sources are properly identified in the future, the residual noise level is around two orders of magnitude lower than the LISA sensitivity curve, thus having a limited effect on the data analysis (see the Bottom Panel). Considering the abovementioned points, we highlight the importance of searching and identifying bursting GW sources in LISA detection. Particularly, we may need to develop portable GW templates for these sources in the matched filtering or adopt other time-frequency approaches, such as power stacking [1] and the Q-transform [98, 99], to characterize the GW bursts from astrophysical sources in LISA data analysis.
Additionally, in realistic detection, the antenna pattern of LISA undergoes rotation across the sky, causing its response to gravitational waves to change with time. As the bursting BBHs in the Milky Way predominantly contribute to the stochastic background and may have an anisotropic distribution, the amplitude of the GW burst background can vary considerably throughout the year. This observation is consistent with the pattern illustrated in Figure 1 [see, e.g., 153, for a similar case].
@@ Acknowledgements.
We thank the anonymous referee for useful comments. ZX acknowledges partial support from the Bhaumik Institute for Theoretical Physics summer fellowship. ZX, SN, and EM acknowledge the partial support from NASA ATP 80NSSC20K0505 and from NSF-AST 2206428 grant and thank Howard and Astrid Preston for their generous support. This work was supported by the Science and Technology Facilities Council Grant Number ST/W000903/1 (to BK).